# Fine-tuning Multimodal Transformers on Edge: A Parallel Split Learning Approach

### v1

@@ 4.1 MPSL Evaluation
We evaluate MPSL on $7$ datasets and compare its performance against two distributed baselines in Tables 1 and 2. As shown in Table 1, in classification tasks,  MPSL consistently matches FedAvg across all modality pairs and tasks while significantly reducing client-side computation. By offloading the computationally intensive model components to the server, MPSL achieves over $250$-fold reduction in GFLOPs requiring only $1.0$M trainable parameters – a 97.71% decrease compared to FedAvg. Notably, MPSL outperforms FedAvg in certain settings, such as MELD, demonstrating its potential for both efficiency and performance. Alternatively, in image-text retrieval, as shown in Table 2, all $3$ distributed ML approaches exhibit a notable performance drop compared to their centralized counterparts. Despite this, MPSL outperforms both FL baselines while maintaining client-side computational efficiency similar to that observed in classification tasks. We further elaborate on the performance gap relative to centralized settings for image-text retrieval tasks in Section 4.2.1.
In terms of communication overhead, MPSL relies solely on the size of communicated smashed data (i.e., activations), rather than the number of trainable parameters, as in most FL regimes. As depicted in Figure 3, while FedAvg demonstrates greater communication efficiency for smaller ViT encoder depths (e.g., ViT-Tiny with $6$ blocks and ViT-Small with $8$ blocks), this advantage diminishes with larger encoder architectures, which are common in multimodal transformers Girdhar et al. (2023). Moreover, FedCLIP – which utilizes lightweight adapters atop a frozen backbone – achieves the lowest communication cost in our evaluations; however, its performance remains limited, with MPSL consistently outperforming it across all tasks while requiring significantly fewer GFLOPs. Consequently, MPSL emerges as a promising alternative for both computation and communication-efficient decentralized multimodal learning, particularly for resource-constrained edge devices.
@@ 4.2.1 Batch size effect
Figure 4: Visualization of the effect of batch size on embedding spaces after PCA for Flickr30K and MS-COCO with $N$=$25$. Blue represents embeddings from centralized models, while green and red indicate MPSL’s embeddings with batch size of $50$ and $200$, respectively.
@@ 4.2.2 Trainable parameters
Figure 6: Impact of encoder depth on model performance for COCO-QA.
@@
Table 4: Performance across tasks when fine-tuning the last 1, 6, or all ViT blocks, with 25 and 100 clients. Results report test accuracy and top-1 Image-to-Text recall (for Flickr30K and MS-COCO), normalized to centralized baselines.
@@ 4.2.3 Fusion Type
Table 5: Impact of early vs. late modality fusion on model performance across classification tasks. We report test set accuracy. V: Vision, A: Audio, T: Text.
@@ 5 Conclusions
We introduce MPSL, a Multimodal Parallel Split Learning framework that can be used to fine-tune multimodal transformers in a distributed manner with lightweight client computational burden. It uses an aggregated loss function of local client losses and a single unified backward pass on the server to eliminate the need for label sharing, synchronization of client-side models, and maintaining per-client sub-models on the server, alongside employing lightweight modality-specific tokenizers on the client-side to reduce computational overhead. Our experiments across multiple datasets, types of tasks, and pairs of modalities demonstrate MPSL’s effectiveness in fine-tuning such models with comparable or superior performance to FL regimes while significantly decreasing client-side computational burden, and achieving superior scalability in communication cost with model growth.

### v2

@@ 4.1 MPSL Evaluation
Model Performance. We evaluate MPSL on $7$ datasets and compare its performance against two distributed baselines in Tables 1 and 2. As shown in Table 1, in classification tasks, MPSL consistently matches FedAvg across all modality pairs and tasks while significantly reducing client-side computation. By offloading the computationally intensive model components to the server, MPSL achieves over a $250$-fold reduction in GFLOPs, requiring only $1.0$M trainable parameters – a 97.71% decrease compared to FedAvg. Notably, MPSL outperforms FedAvg in certain settings, such as MELD, demonstrating its potential for both efficiency and performance. In contrast, for image-text retrieval (Table 2), all $3$ distributed approaches exhibit a notable performance drop compared to centralized training. Despite this, MPSL surpasses both FL baselines while maintaining client-side efficiency on par with classification tasks. We further elaborate on the gap between distributed and centralized performance for retrieval in Section 4.2.1.
Convergence Analysis. We compare the convergence speed of MPSL and FedAvg (excluding FedCLIP due to its poor performance; see Table 1) by measuring the number of rounds to reach 95% of peak accuracy. As shown in Figure 2MPSL converges significantly faster, especially in the $N{=}100$ regime where data is highly distributed. This stems from their differing update strategies: FedAvg relies on local backpropagation and global averaging, which can lead to client drift, while MPSL uses a single server-side backpropagation over aggregated losses, providing a stabilizing global signal. This enables MPSL to converge more quickly and remain stable even in distributed multimodal settings, where modality misalignment compounds the effects of data heterogeneity and further hinders optimization.
Figure 2: Average number of rounds to reach 95% of peak performance. Numbers above bars denote average final accuracy.
@@
Communication Cost. Unlike FL regimes that depend on transmitting trainable parameters, MPSL only communicates smashed data (i.e., activations), decoupling communication from model size. As shown in Figure 3, FedAvg can be more communication-efficient for small ViT models (e.g., Tiny, Small), but this benefit fades with larger encoders — common in multimodal transformers Girdhar et al. (2023). FedCLIP, using lightweight adapters on a frozen backbone, achieves the lowest communication cost overall; however, its performance remains limited (see Table 1). In contrast, MPSL consistently outperforms FedCLIP across all tasks while maintaining lower computation. This highlights MPSL’s suitability for scalable multimodal learning under strict communication constraints.
@@ 4.2.1 Batch size effect
Figure 4: Effect of batch size on embedding spaces after PCA for Flickr30K and MS-COCO with $N$=$25$. Blue represents embeddings from centralized models, while green and red indicate MPSL’s embeddings with batch size of $50$ and $200$, respectively.
@@ 4.2.2 Trainable parameters
Figure 6: Impact of encoder depth on performance for COCO-QA.
@@
Table 4: Performance across tasks when fine-tuning the last 1, 6, or all ViT blocks, with 25 and 100 clients. We report test accuracy and top-1 Image-to-Text recall (for Flickr30K and MS-COCO), normalized to centralized baselines.
@@ 4.2.3 Fusion Type
Table 5: Impact of early vs. late modality fusion on performance across classification tasks. V: Vision, A: Audio, T: Text.
@@ 5 Conclusions
We introduce MPSL, a Multimodal Parallel Split Learning framework that can be used to fine-tune multimodal transformers in a distributed manner with lightweight client computational burden. It uses an aggregated loss function of local client losses and a single unified backward pass on the server to eliminate the need for label sharing, synchronization of client-side models, and maintaining per-client sub-models on the server, alongside employing lightweight modality-specific tokenizers on the client-side to reduce computational overhead. Our experiments across multiple datasets, types of tasks, and pairs of modalities demonstrate MPSL’s effectiveness in fine-tuning such models with comparable or superior performance to FL regimes while significantly decreasing client-side computational burden, and achieving superior scalability in communication cost with model growth. Future work involves addressing real-world challenges such as heterogeneous edge hardware, client dropouts, and missing modalities across clients.
Acknowledgments
This work was funded by the DAIS project, which has received funding from ECSEL Joint Undertaking under grant agreement No 101007273, and the Bio-Curity project, which has received funding from Eureka Cluster Xecs under grant agreement No 2022016.